OpenStudy (bbb911):

A square has an area of 49 square inches. If the same amount is added to the length and removed from the width, the resulting rectangle has an area of 45 square inches. Find the dimensions of the rectangle.

OpenStudy (ash2326):

@bbb911 Do you know the relation between area of a square and its side?

OpenStudy (bbb911):

not really good with that

OpenStudy (campbell_st):

ok... so you know that \[x^2 = 49\] so the side length is 7 so let y be the amount added and subtracted and you get ( 7 - y)(7 + y) = 45 so \[49 - y^2 = 45\] just solve for y... and only use the positive value

OpenStudy (ash2326):

Suppose a a square is having side as "x" We know that all 4 sides of a square are equal.  Area is length times breadth for a square length=breadth=x So Area will be? can you tell?

OpenStudy (bbb911):

Okay one sec while i try and solve it :)

OpenStudy (ash2326):

Sure!!

OpenStudy (bbb911):

is it 5 in by 9 in

OpenStudy (campbell_st):

thats correct...

OpenStudy (bbb911):

yay!!!! :D

OpenStudy (bbb911):

Thanks you too :D